OpenStudy (mollykenney):

Let f (x) -2x + 5 and g(x) x^3 What is (g - f )(x)? A. x^3 - 2x + 5 B. x^3 + 2x - 5 C. -x^3 - 2x + 5 D. -x^3 + 2x - 5

jimthompson5910 (jim_thompson5910):

are the two functions f(x) = -2x+5 and g(x) = x^3 right?

OpenStudy (mollykenney):

yes

jimthompson5910 (jim_thompson5910):

(g-f)(x) is the same as saying g(x) - f(x)

jimthompson5910 (jim_thompson5910):

g(x) - f(x) = x^3 - (-2x+5) g(x) - f(x) = x^3 + 2x - 5

jimthompson5910 (jim_thompson5910):

be sure to subtract ALL of f(x) from g(x) so distribute that negative through to each term in f(x)

OpenStudy (mollykenney):

thank you so much @jim_thompson5910 i understand it now

jimthompson5910 (jim_thompson5910):

you're welcome